# Efficient and Generalizable Certified Unlearning: A Hessian-free Recollection Approach

### v3

@@ Abstract
Machine unlearning strives to uphold the data owners’ right to be forgotten by enabling models to selectively forget specific data. Recent advances suggest precomputing and storing statistics extracted from second-order information and implementing unlearning through Newton-style updates. However, the theoretical analysis of these works often depends on restrictive assumptions of convexity and smoothness, and those mentioned operations on Hessian matrix are extremely costly. As a result, applying these works to high-dimensional models becomes challenging. In this paper, we propose an efficient Hessian-free certified unlearning. We propose to maintain a statistical vector for each data, computed through affine stochastic recursion approximation of the difference between retrained and learned models. Our analysis does not involve inverting Hessian and thus can be extended to non-convex non-smooth objectives. Under same assumptions, we demonstrate advancements of proposed method beyond the state-of-the-art theoretical studies, in terms of generalization, unlearning guarantee, deletion capacity, and computation/storage complexity, and we show that the unlearned model of our proposed approach is close to or same as the retrained model. Based on the strategy of recollecting statistics for forgetting data, we develop an algorithm that achieves near-instantaneous unlearning as it only requires a vector addition operation. Experiments demonstrate that the proposed scheme surpasses existing results by orders of magnitude in terms of time/storage costs, while also enhancing accuracy.
@@ 1 Introduction
A straightforward unlearning algorithm is to retrain a learned model from scratch, which is however impractical as it comes with unaffordable monetary and time costs. It is especially challenging in time-sensitive applications such as fraud detection or online learning, as neglecting timeliness may have a detrimental impact on system performance and responsiveness. Moreover, attaining a model entirely identical to the one achieved through retraining for an unlearning method represents a notably demanding criterion. Therefore, many approximate methods have been proposed aiming to minimize the impact of forgetting data. Inspired by differential privacy, [1] introduced a relaxed definition of certified unlearning, which aims to make unlearned model indistinguishable from retrained model.
@@
Second-order unlearning methods have been recently studied due to their rigorous unlearning certification and the generalization guarantee. Sekhari et al. [2] initially proposed Newton update method and introduced the definition of deletion capacity to understand the relationship between generalization performance and the amount of deleted samples. These method and concepts were later extended, e.g. [3, 4, 5, 6]. The key idea of these works is to extract Hessian information of the loss function and achieve computational and memory efficiency by pre-computing and storing these second-order statistics prior to any deletion request. In particular, Sekhari et al. [2] and Suriyakumar et al. [3] studies reveal that for models that approximate the minimization of convex and sufficiently smooth empirical minimization objectives, unlearning can be efficiently executed through a Newton style step, facilitated by pre-storing Hessian matrix. While these works have demonstrated the effectiveness of second-order method to approximate retrained model with limited updates, they do have a few fundamental limitations. First, most existing certified unlearning algorithms require the explicit pre-storage of the Hessian matrix (or its inverse) and are hence expensive to implement, especially in high-dimensional models. This forms a stark contrast to many algorithms leveraging Hessian-free solvers, e.g. exploit the Hessian vector product technique [7] without explicitly computing Hessian matrix. Second, their analyses are not applicable to most deep neural networks due to their requirement of convexity assumption. We are thus motivated to answer the following key question:
How to realize certified unlearning without directly computing Hessian and restrictive assumptions?
@@
In contrast to prior works, this paper aims to study the efficient second-order unlearning methods from a new Hessian-free perspective. Specifically, we analyze the difference between the learned and retrained models for each sample by recollecting the trajectory discrepancy. Subsequently, we demonstrate that these discrepancies can be approximated and characterized through an affine stochastic recursion. We summarize our main contributions related to theoretical and experimental results below, and the contributions related to algorithm efficiency in Table 1.
@@
    Hessian-Free Analysis. We propose a method to extract Hessian information by analyzing each sample’s impact on the training trajectory without explicitly computing the Hessian. Our method allows us to leverage the Hessian-free technique, which enjoys better scalability compared to existing second-order unlearning methods. Since our analysis avoids inverting Hessian, it can also be extended to non-convex non-smooth objectives. To our best knowledge, our work is the first to establish non-convex theoretical analysis of certified unlearning. (Section 3).
@@
    Improved Guarantees and Complexity. We demonstrate advancements of proposed method compared to the state-of-the-art theoretical studies, in terms of generalization (See Theorem Theorem 5), unlearning guarantee (See Theorem 4), deletion capacity (See Theorem 8), and computation/storage complexity (See Subsection 4.4 and Table 1). Especially in Theorem 4, we show the proposed method can almost achieve the same model as the retrained model. (Section 4).
@@
    Instantaneous Unlearning Algorithm. Building upon the aforementioned analysis, we develop a Hessian-free unlearning algorithm. To our best knowledge, our algorithm is near-instantaneous and currently the most efficient as it only requires a vector addition operation. (Section 4).
@@
    Comprehensive Evaluation. We conduct experimental evaluations using a wider range of metrics compared to previous theoretical studies for certified unlearning. The experimental results verify our theoretical analysis and demonstrate that our proposed approach surpasses previous certified unlearning works. In particular, our algorithm incurs the millisecond-level unlearning runtime to forget per sample with minimal performance degradation. (Section 5).
Table 1: $d$ is the dimensionality of the model parameters, and $n$ is the total size of the dataset.
Methods Pre-computation time Storage Unlearning time (per sample)
Sekhari et al. [2] $\mathcal{O}(nd^{2})$ $\mathcal{O}(d^{2})$ $\mathcal{O}(d^{3})$
Suriyakumar et al. [3] $\mathcal{O}(d^{3}+nd^{2})$ $\mathcal{O}(d^{2})$ $\mathcal{O}(d^{2})$
Proposed Hessian-free method $\mathcal{O}(nd)$ $\mathcal{O}(nd)$ $\mathcal{O}(d)$
@@ 1.1 Related Work
Machine unlearning can be traced back to [8] and their work defines the notion of completeness, demanding that the models obtained through retraining and unlearning exhibit complete consistency. Exact Unlearning methods, e.g., [9, 10, 11, 12, 13, 14], train submodels based on different dataset partitions and aggregation strategies (See detailed exact unlearning background in Appendix A). Approximate Unlearning, e.g. [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25] seeks to minimize the impact of forgetting data to an acceptable level to tradeoff for computational efficiency, reduced storage costs, and flexibility (See detailed approximate unlearning background in Appendix A). In particular, [1] introduced a relaxed $(\epsilon,\delta)$-certified unlearning definition, in anticipation of ensuring the output distribution of the unlearning algorithm is indistinguishable from that of the retraining. Building upon this framework, diverse methodologies, e.g. [2, 3, 4, 6, 26, 27], have been devised.
In the certified unlearning definition, the Hessian-based approaches [2, 3, 4] are aligned with our methodology, which involves pre-computing data statistics that extract second-order information from the Hessian matrix, and utilizing these precomputed ‘recollections’ to facilitate data forgetting. Specifically, (1) Newton Step (NS). [2] proposes, through the precomputation and storage of the Hessian offline, to achieve forgetting upon the arrival of a forgetting request using a Newton step. However, this method necessitates strong assumptions, e.g. the objective function is required to be strongly convex. Its unlearning time, involving $\mathcal{O}(d^{3})$, becomes nearly impractical for over-parameterized deep neural networks. (2) Infinitesimal Jacknife (IJ). [3, 4] extends the work of NS, attaining strong convexity through a non-smooth regularizer by leveraging the proximal IJ with reduced unlearning time $\mathcal{O}(d^{2})$. However, the basic issues still exist, i.e., existing certified unlearning methods, especially those focused on establishing theoretical results, have largely focused on convex and smooth settings while requiring expensive Hessian matrix computations. On the contrary, our method provides a breakthrough for certified unlearning in a non-convex non-smooth setting. In addition, as indicated in Table 1, our method breaks the curse of model parameters dimensionality. Moreover, prior works necessitate maintaining the Hessian (or its inverse) to address the ongoing deletion requests, which poses the risk of violating the right to be forgotten and entails user privacy. Instead, our method goes beyond ‘recollecting to forget’ by also ‘forgetting the recollections’.
@@ 2 Problem Formulation
Learning   Let $\ell(\mathbf{w};z)$ be the loss function for a given parameter $\mathbf{w}\!\in\!\mathbb{R}^{d}\!$ and on instance $z$ over an instance space $\mathcal{Z}$. Learning can be expressed as the population risk minimization problem:
@@
A standard approach to solve the problem (2) is the stochastic gradient descent (SGD), iterating towards the direction with stepsize $\eta$ of the negative gradient (stochastically):
@@
for the index of epoch update $e\in\{0,...,E\}$ and the index of batch update $b\in\{0,...,B\}$, where $E$ is total epochs and $B$ is batches per epoch. Here, $\mathcal{B}_{e,b}$ denotes the dataset that contains the data sampled in the $e$-th epoch’s $b$-th update, and $|\mathcal{B}_{e,b}|$ is the size of $\mathcal{B}_{e,b}$, where $|\mathcal{B}|$ is the maximum batchsize.
Unlearning   Consider a scenario where a user requests to remove a forgetting dataset $U\!\!=\!{\{u_{j}\}}_{j=1}^{m}\!\subseteq\!\mathcal{S}$. To unlearn $U$, retraining from scratch is considered a naive unlearning method, which aims to minimize $F_{\mathcal{S}\backslash U}(\mathbf{w})$. Specifically, since each sample takes place only in one (mini-)batch per epoch, we define the data point $u_{j}$ in $U$ to be sampled in $\mathcal{B}_{e,b(u_{j})}$ during the $e$-th epoch’s $b(u_{j})$-th batch update in the learning process, where $b(u_{j})$ represents the batch update index when $u_{j}$ is sampled. Due to the removal of sample $u_{j}$, the batch $\mathcal{B}_{e,b(u_{j})}$ is reduced to $\mathcal{B}_{e,b(u_{j})}\backslash\{u_{j}\}$. In accordance with the linear scaling rule discussed in [28], i.e., in order to effectively leverage batch sizes, one has to adjust the stepsize in tandem with the batch size. The retraining stepsize thus is concurrently scaled to $\eta_{e,b(u_{j})}\frac{|\mathcal{B}_{e,b(u_{j})}\backslash\{u_{j}\}|}{|\mathcal{B}_{e,b(u_{j})}|}$. Therefore, the retraining update rule in $e$-th epoch is given by,
@@
Specifically, for the data point $u_{j}$ sampled into $\mathcal{B}_{e,b}$ during $e$-th epoch’s $b(u_{j})$-th update, we have:
@@
To avoid the unaffordable costs incurred by the retraining algorithm, we aim to approximate the retrained model through limited updates on the learned model $\mathbf{w}_{e,b+1}$. We thus consider the relaxed definition of model indistinguishability defined by Sekhari et al. [2].
@@ 3 Main Intuition and Solution Justification
Before detailing our analysis, let us informally motivate our main method. Specifically, we focus on studying the discrepancy between the learning and the retraining model by analyzing the impact of each sample on the update trajectory. Our main observation is that the difference between the retraining and learning models analyzed through SGD recursion, can be described by an affine stochastic recursion. See full proofs of Section 3 in Subsection B.1 and Subsection B.2.
Unlearning a single data sample. We start with designing an approximator for a single sample (i.e., $U\!=\!\{u\}$) on the training updates (the learning update and retraining update), by considering the difference between the learning model in (3) and the retraining model obtained without the knowledge of $u$ in (4). In particular, we consider the following two cases of model updates in the $e$-th epoch.
@@
From the Taylor expansion of $\nabla\ell(\mathbf{w}^{-u}_{e,b};z_{i})$ around $\mathbf{w}_{e,b}$, we have $\nabla\ell(\mathbf{w}_{e,b}^{-u};z_{i})=\nabla\ell(\mathbf{w}_{e,b};z_{i})+\nabla^{2}\ell(\mathbf{w}_{e,b};z_{i})(\mathbf{w}_{e,b}^{-u}-\mathbf{w}_{e,b})$$+{o(\mathbf{w}_{e,b}^{-u}-\mathbf{w}_{e,b})}.$ Let $\mathbf{H}_{e,b}\!\!=\!\!\sum_{i\in\mathcal{B}_{e,b}}\!\!\!\!\nabla^{2}\ell(\mathbf{w}_{e,b};z_{i})$ denote the Hessian of the loss during the $e$-th epoch’s $b$-th update, we can then approximate the SGD recursion (7) as
$\mathbf{w}_{e,b+1}^{-u}-\mathbf{w}_{e,b+1}\approx\Big{(}\mathbf{I}-\frac{\eta_{e,b}}{|\mathcal{B}_{e,b}|}\mathbf{H}_{e,b}\Big{)}(\mathbf{w}_{e,b}^{-u}-\mathbf{w}_{e,b}).$ (8)

### v4

# Hessian-Free Online Certified Unlearning

@@ Abstract
Machine unlearning strives to uphold the data owners’ right to be forgotten by enabling models to selectively forget specific data. Recent advances suggest pre-computing and storing statistics extracted from second-order information and implementing unlearning through Newton-style updates. However, the Hessian matrix operations are extremely costly and previous works conduct unlearning for empirical risk minimizer with the convexity assumption, precluding their applicability to high-dimensional over-parameterized models and the nonconvergence condition. In this paper, we propose an efficient Hessian-free unlearning approach. The key idea is to maintain a statistical vector for each training data, computed through affine stochastic recursion of the difference between the retrained and learned models. We prove that our proposed method outperforms the state-of-the-art methods in terms of the unlearning and generalization guarantees, the deletion capacity, and the time/storage complexity, under the same regularity conditions. Through the strategy of recollecting statistics for removing data, we develop an online unlearning algorithm that achieves near-instantaneous data removal, as it requires only vector addition. Experiments demonstrate that our proposed scheme surpasses existing results by orders of magnitude in terms of time/storage costs with millisecond-level unlearning execution, while also enhancing test accuracy.
footnotemark: footnotetext: Corresponding Author
@@ 1 Introduction
A straightforward unlearning algorithm is to retrain a learned model from scratch, which is however impractical as it incurs unaffordable costs. It is especially challenging in time-sensitive applications such as fraud detection or online learning, as delayed updates upon receiving a deletion request would impede the system’s ability to respond effectively to subsequent inputs, compromising its performance and reliability. Moreover, attaining a model entirely identical to the one achieved through retraining for an unlearning method represents a notably demanding criterion. Therefore, many approximate methods have been proposed aiming to minimize the impact of forgetting data. Inspired by differential privacy, Guo et al. (2020) introduced a relaxed definition of certified unlearning, which aims to make the output distribution of unlearned algorithm indistinguishable from that of the retraining.
@@
Second-order unlearning methods have been recently studied due to their rigorous unlearning certification and the generalization guarantee. Sekhari et al. (2021) initially proposed the Newton update method and introduced the definition of deletion capacity to understand the relationship between generalization performance and the amount of deleted samples. These method and concepts were later extended, e.g. Suriyakumar et al. (2022); Liu et al. (2023); Mahadevan & Mathioudakis (2021); Chien et al. (2022). The key idea of these works is to extract Hessian information of the loss function and achieve computational and memory efficiency by pre-computing and storing these second-order statistics prior to any deletion request. In particular, Sekhari et al. (2021) and Suriyakumar et al. (2022) revealed that for models that approximate the minimization of convex and sufficiently smooth empirical minimization objectives, unlearning can be efficiently executed through a Newton style step, facilitated by pre-storing Hessian matrix. These methods and theoretical analyses are subsequently extended to graph learning (Chien et al. (2022)) and minimax learning (Liu et al. (2023)) scenarios.
While these works have demonstrated the effectiveness of the second-order method to approximate retrained models with limited updates, they do have a few fundamental challenges: existing second-order algorithms (i) lack the capability to handle multiple deletion requests in online because they require explicit precomputing Hessian (or its inverse), and (ii) the learned model needs to be empirical risk minimizer, (iii) with assumption of convexity to ensure the invertibility of Hessian. Consequently, these limitations preclude their applicability to high-dimensional models (e.g., over-parameterized neural networks) for computing the Hessian. For non-convex problems (e.g., deep networks), we cannot guarantee the invertibility of the Hessian, and it is non-trivial for a learned model to converge to an empirical risk minimizer. Even for convex problems, the non-stationarity of data (Hazan (2016)) is another common issue in the incremental model update scenarios (e.g., continual learning) where training is sequential but not concurrent, making it challenging for the model to adapt and converge to the optimal solution. Consequently, we are motivated to answer the following key question:
(Q) How can efficient second-order unlearning be achieved without incurring the high cost of Hessian operations, and how can effective certified unlearning be implemented under non-convex assumptions for the objectives and non-convergence conditions for the models?
@@
In this paper, we take the first step towards tackling (Q) from an innovative Hessian-free perspective. In contrast to existing schemes that extract second-order information exclusively from the optimal learned model, we leverage details from learning models during training procedure. We analyze the difference between the learning model and the retraining model by recollecting the model update trajectory discrepancy. Subsequently, we demonstrate that the discrepancies can be characterized through an affine stochastic recursion. We summarize our main contributions below and in Table 1.
@@
    Hessian-Free Model Update Method. We propose a method to extract the Hessian information by analyzing each sample’s impact on the training trajectory without explicitly computing the Hessian matrix. Our Hessian-free method enables efficient handling of deletion requests in an online manner. Moreover, the proposed method does not require that the learned model is empirical risk minimizer, which enjoys better scalability compared to existing second-order methods. (Section 3).
@@
    Improved Guarantees and Complexity. Under identical regularity conditions, we analytically show that our proposed method outperforms the state-of-the-art theoretical studies in terms of unlearning guarantee (See Corollary 5), generalization guarantee (See Theorem 6), deletion capacity (See Theorem 7), and computation/storage complexity (See Subsection 4.4). Moreover, the unlearning guarantee can be applied to both convex and non-convex settings. (Section 4).
@@
    Instantaneous Unlearning Algorithm. Building upon the aforementioned analysis, we develop a Hessian-free certified unlearning algorithm. To our knowledge, our algorithm is near-instantaneous and is currently the most efficient as it requires only a vector addition operation. (Section 4).
@@
[…]
Table 1: Summary of Results. Here, $d$, $n$, and $m$ denote the model parameters size, training dataset size, and forgetting dataset size, respectively. The unlearning guarantee is derived under the same regularity conditions as prior second-order unlearning studies, with $\rho$ being a constant less than 1.
Methods Computation time Storage Unlearning time Unlearning guarantee
Sekhari et al. (2021) $\mathcal{O}(nd^{2})$ $\mathcal{O}(d^{2})$ $\mathcal{O}(d^{3}+md^{2}+md)$ $\mathcal{O}(m^{2}/n^{2})$
Suriyakumar et al. (2022) $\mathcal{O}(d^{3}+nd^{2})$ $\mathcal{O}(d^{2})$ $\mathcal{O}(d^{2}+md)$ $\mathcal{O}(m^{2}/n^{2})$
Proposed method $\mathcal{O}(n^{2}d)$ $\mathcal{O}(nd)$ ${\mathcal{O}(md)}$ $\mathcal{O}(m\rho^{n})$
@@ 1.1 Related Work
Machine unlearning can be traced back to Cao & Yang (2015), which defines "the completeness of forgetting", demanding the models obtained through retraining and unlearning exhibit consistency. Exact Unlearning, e.g., Bourtoule et al. (2021); Brophy & Lowd (2021); Schelter et al. (2021); Yan et al. (2022); Chen et al. (2022b; a), typically train the submodels based on different dataset partitions and aggregation strategies (See elaborated exact unlearning background in Appendix A). Approximate Unlearning, e.g. Wu et al. (2020); Nguyen et al. (2020); Golatkar et al. (2020a); Izzo et al. (2021); Mehta et al. (2022); Wu et al. (2022); Tanno et al. (2022); Becker & Liebig (2022); Jagielski et al. (2023); Warnecke et al. (2023); Tarun et al. (2023), seeks to minimize the impact of forgetting data to an acceptable level to tradeoff for computational efficiency, reduced storage costs, and flexibility (See elaborated approximate unlearning background in Appendix A). In particular, Guo et al. (2020) introduces a relaxed definition of $(\epsilon,\delta)$-unlearning, in anticipation of ensuring the output distribution of an unlearning algorithm is indistinguishable from that of the retraining. Building upon this framework, diverse methodologies, e.g. Sekhari et al. (2021); Suriyakumar et al. (2022); Gupta et al. (2021); Neel et al. (2021); Liu et al. (2023); Chien et al. (2022), have been devised.
In the certified unlearning definition, the Hessian-based approaches (Sekhari et al. (2021); Suriyakumar et al. (2022); Liu et al. (2023)) are aligned with our methodology, which involves pre-computing data statistics that extract second-order information from the Hessian and using these precomputed "recollections" to facilitate data forgetting. Specifically, (1) Newton Step (NS). Sekhari et al. (2021) proposed, through the precomputation and storage of the Hessian offline, to achieve forgetting upon the arrival of a forgetting request using a Newton step. However, NS is limited by its high computational complexity of $\mathcal{O}(d^{3})$, making it impractical for large-scale deep models. Furthermore, it can only perform certified unlearning for empirical minimizers under convexity assumption. (2) Infinitesimal Jackknife (IJ). Suriyakumar et al. (2022) extended the work of NS, attaining strong convexity through a non-smooth regularizer by leveraging the proximal IJ with reduced unlearning time $\mathcal{O}(d^{2})$. However, a few important challenges still remain, i.e., existing second-order methods still require expensive Hessian computations and perform forgetting for empirical risk minimizer with the convexity assumption. In contrast, our method overcomes the limitations of existing second-order methods by achieving near-instantaneous unlearning and performing forgetting for models in any training step. Hence, our approach is applicable to scenarios where models are continuously updated.
@@ 2 Problem Formulation
Learning. Let $\ell(\mathbf{w};\!z)$ be loss function for a given parameter $\mathbf{w}$ over parameter space $\mathcal{W}$ and sample $z$ over instance space $\mathcal{Z}$. Learning can be expressed as a population risk minimization problem:
@@
A standard approach to solve the problem (2) is the stochastic gradient descent (SGD), iterating towards the direction with stepsize $\eta_{e,b}$ of the negative gradient (stochastically):
@@
for the index of epoch update $e\in\{0,...,E\!+\!1\}$ and batch update $b\in\{0,...,B\!+\!1\}$, where $E$ is total epochs and $B$ is batches per epoch. Here, $\mathcal{B}_{e,b}$ denotes the set that contains the data sampled in the $e$-th epoch’s $b$-th update, and $|\mathcal{B}_{e,b}|$ is the size of $\mathcal{B}_{e,b}$, where $|\mathcal{B}|$ is the maximum batchsize.
Unlearning. Consider a scenario where a user requests to remove a forgetting dataset $U\!\!=\!{\{u_{j}\}}_{j=1}^{m}\!\subseteq\!\mathcal{S}$. To unlearn $U$, retraining from scratch is considered a naive unlearning method, which aims to minimize $F_{\mathcal{S}\backslash U}(\mathbf{w})$. Specifically, for ease of understanding, each sample appears only in one (mini-)batch per epoch, we define the data point $u_{j}$ in $U$ to be sampled in $\mathcal{B}_{e,b(u_{j})}$ during the $e$-th epoch’s $b(u_{j})$-th update in the learning process, where $b(u_{j})$ represents the batch update index when $u_{j}$ is sampled. In accordance with the linear scaling rule discussed in Goyal et al. (2017), i.e. in order to effectively leverage batch sizes, one has to adjust the step size in tandem with batch size. In other words, the step size to batch size ratio during retraining is consistent with the learning process. Therefore, the update rule of retraining process during $e$-th epoch is given by,
@@
Specifically, for the training data $u_{j}$ sampled into $\mathcal{B}_{e,b(u_{j})}$ in the learning process, the removal of sample $u_{j}$ in the retraining process results in the following update:
@@
To avoid the unaffordable costs incurred by the retraining algorithm, we aim to approximate the retrained model through limited updates on the learned model $\mathbf{w}_{e,b+1}$. We thus consider the relaxed definition of model indistinguishability defined by Sekhari et al. (2021).
@@
Although classical Gaussian mechanisms (Dwork & Roth (2014)) assume $0<\epsilon\leq 1$, Balle & Wang (2018) addressed $\epsilon>1$ by using Gaussian cumulative density function instead of a tail bound approximation. For consistency with prior works, we use classical mechanisms in the main text.
@@ 3 Main Intuition and Solution Justification
Before detailing our analysis, let us informally motivate our main method. Specifically, we focus on studying the discrepancy between the learning and the retraining model by analyzing the impact of each sample on the update trajectory. Our main observation is that the difference between the retraining and learning models analyzed through SGD recursion can be described by an affine stochastic recursion. Suppose retraining and learning processes share the same initialization, such as using the same random seed or pre-trained model. We now commence the main procedure as follows:
Unlearning a single data sample. We start with designing an approximator for a single sample (i.e., $U\!=\!\{u\}$) on the training (the learning and retraining) updates, considering the difference between the learning model in Eq. 3 and the retraining model obtained without the knowledge of $u$ in Eq. 4. In particular, we consider the following two cases of model updates in the $e$-th epoch.
@@
Suppose loss function $\ell\in G^{2}\left(\mathbb{R}^{d}\right)$, i.e. $\ell$ is twice continuously differentiable. Therefore, from the Taylor expansion of $\nabla\ell(\mathbf{w}^{-u}_{e,b};z_{i})$ around $\mathbf{w}_{e,b}$, we have $\nabla\ell(\mathbf{w}_{e,b}^{-u};z_{i})\approx\nabla\ell(\mathbf{w}_{e,b};z_{i})+\nabla^{2}\ell(\mathbf{w}_{e,b};z_{i})(\mathbf{w}_{e,b}^{-u}-\mathbf{w}_{e,b})$. Let $\mathbf{H}_{e,b}\!\!=\!\!{1}/{|\mathcal{B}_{e,b}|}\sum_{i\in\mathcal{B}_{e,b}}\!\!\!\!\nabla^{2}\ell(\mathbf{w}_{e,b};z_{i})$ denote the Hessian matrix of the loss function, we can then approximate the recursion of Eq. 7 as
$\mathbf{w}_{e,b+1}^{-u}-\mathbf{w}_{e,b+1}\approx(\mathbf{I}-{\eta_{e,b}}\mathbf{H}_{e,b})(\mathbf{w}_{e,b}^{-u}-\mathbf{w}_{e,b}).$ (8)